OpenStudy (anonymous):

find three arithmetic means between - 2 and 12 would it be 8 10 12

OpenStudy (anonymous):

12 = -2 + 4d 14 = 4d 3.5 = d -2 + 3.5 = 1.5 1.5 + 3.5 = 5 5 + 3.5 = 8.5 1.5, 5, and 8.5 are the three arithmetic means.

OpenStudy (anonymous):

Do you understand?

OpenStudy (anonymous):

im a little confused

OpenStudy (anonymous):

Where are you confused?

OpenStudy (anonymous):

how did you do it like i get how you added them im sorry i really dont know this tuff

OpenStudy (anonymous):

It's okay so to find the three arithmetic means, you need to subtract the two numbers. 12 - (-2) = 14 and then divide the difference by how many means you want plus one, since it is three, you need to divide it by four; 14/4 then add this factor to your lower limit -2 + 14/4 -8/4 + 14/4 (getting the lcd) = 6/4 or 3/2 (first arithmetic mean) for your second arithmetic mean, just add the factor to your first arithmetic mean 6/4 + 14/4 = 20/4 or 5 (second arithmetic mean) 20/4 + 14/4 = 34/4 or 17/2 (third arithmetic mean) to check if you were able to got the right answers, add the factor to your last arithmetic and the sum should be your upper boundary. 34/4 + 14/4 = 48/4 or 12

OpenStudy (anonymous):

Hope that helps!

OpenStudy (anonymous):

thank you so much that really helps thanks for such a helpful answer i really really appreciate it :)

OpenStudy (anonymous):

welcome!

OpenStudy (anonymous):

:D